Qd= 1000-10P Qs= 10P

1. If a 20$ subsidy is placed on the Good What is new Pc, Pp, and Q.

2. What is Consumer surplus?

3. What is producer surplus?

4. What is the cost to the government?

5. What is total surplus?

6. What is dead weight loss?
Answer #1

1) New demand is Qd = 1000 - 10(P - 20) or Qd = 1200 - 10P

New supply is Qs = 10P

New equilibrium has

Qs = Qd

10P = 1200 - 10P

P = 1200/20

= $60 per unit (received by sellers)

P = 60 - 20 = $40 (paid by buyers)

Q = 10*60 = 600 units

b) CS = 0.5*(maximum price - price paid)*qty = 0.5*(100 - 40)*600 = $18000

c) PS = 0.5*(price received - minimum price)*qty = 0.5*(60 - 0)*600 = $1800

d) Cost to government = subsidy * quantity = 20*600 = $12000

e) TS = PS + CS = $36000

f) DWL = 0.5*20*100 = $1000